# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v1

@@ VI-B Robustness of the maximum iterations in MBP4+ADOSD4
Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations $T$ for the $[[175,7,7]]$ LCS code.
Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. We note that in ADOSD4, OSD-$w$ is invoked when HRSR fails. As a result, the overall algorithm maintains high accuracy.
However, the failures of the HRSR algorithm can impact the simulation time, as OSD-$w$ is significantly more time-consuming.
Figure 3 illustrates the performance of MBP4+ADOSD4 decoding with varying maximum numbers of iterations, $T=1,2,3,4,5,10,20,50,100,200$. The curve for $T=4$ is already good enough and smooth. While increasing the number of iterations results in slightly improved performance, the curves converge closely around an LER of $2\times 10^{-7}$ to $3\times 10^{-7}$.
In the extreme case of $T=1$, there are few highly reliable bits or valid highly reliable bits, causing the decoder to essentially revert to performing ordinary OSD4 with increased complexity.
We have demonstrated that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. To balance accuracy and efficiency, we typically select $T$ within the range of $50$ to $100$.
VI-C Bivariate bicycle codes
A class of QLDPC codes, referred to as BB codes, was proposed in [55], demonstrating a high error threshold and a significantly higher code rate compared to topological codes. Several attempts at decoding this code family have been presented in [53, 54] under the code capacity noise model. However, the BP+CB decoder in [53] does not perform as well as BP-OSD-0. The BP+GDG decoder in [54] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [54, Figure 4].
For the [[144,12,12]] code, we also include the decoding performance of MBP4+OSD4-2, which is nearly identical to that of MBP4+ADOSD4.
@@ VI-D Generalized hypergraph-product codes
The $[[882,48,16]]$ GHP code [42] is highly-degenerate, which causes conventional BP to perform poorly on this code. With minimum distance of 16, we can gauge the effectiveness of a good decoding algorithm on this GHP code.
The performance curves of different combinations of MBP and OSD are shown in Figure 5, where $\alpha$ is either fixed at 1.6 or chosen by $\alpha(\epsilon)$. It is suggested in [31, Fig. 3] (arXiv version) that $\alpha$ and $\epsilon$ follow the relationship $\alpha(\epsilon)=-0.16\log_{10}(\epsilon)-0.48$, with the coefficients $-0.16$ and $-0.48$ determined through pre-simulations on the [[882, 48]] code. For comparison, we also plot the performance of AMBP4 [31] and the BP+mOSD-($w$, $\lambda$[42].
Figure 5: Comparison of various BP-OSD schemes on the $[[882,48]]$ GHP code. The curve BP+mOSD-(15,15) is from [42].
The results are summarized as follows.
  1. 1.
    MBP4 can be greatly enhanced by OSD.
  2. 2.
    MBP4+OSD4-00 outperforms MBP4+mOSD4-00 (without hard-decision reliability) by roughly half an order of performance. This suggests that the hard-decision history of BP is crucial in determining the reliability order.
  3. 3.
    MBP4+ADOSD4 performs slightly better than MBP4+OSD4-2, while also reducing the time consumption, as shown in Table IV in Section VI-G.
  4. 4.
    Both MBP4+ADOSD4 and MBP+OSD4-2 improve upon BP+mOSD-($15$, $15$[42] by more than an order of magnitude at $\epsilon=0.05$. Note that mOSD-($15$, $15$) essentially enumerates all the 15 less reliable bits.
Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code.
In this paper, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4). Herein we also simulate the performances of MBP4+mOSD based on the reliability metrics of Entropy or Max as shown in Figure 6, including mOSD4  as well. Two conclusions are as follows:
@@
We have conducted additional simulations on a [[625,25,8]] GHP code [43], as shown in Figure 7. In these examples, MBP4+ADOSD4 slightly outperforms both AMBP4 and MBP4+OSD4-2, with all three decoders performing significantly better than BP-OSD-CS60 in [43].
Figure 7: Decoding performance of the [[625,25,8]]] GHP code. Note that the performance curve for BP-OSD-CS60 is taken from [43, Figure 5], with its independent X/Z error rate $p$ rescaled by a factor of $1.5$ to approximate depolarizing error performance.
VI-E 2D topological codes
Next, we consider the decoding performance of our BP+OSD schemes on various 2D topological codes. Since AMBP4 is also capable of decoding various families of topological codes, it serves as a useful reference for evaluating the performance of the BP+OSD approach. It is important to examine the decoding performance when the logical error rate is around $10^{-6}$, as this can reveal whether an iterative decoder exhibits an error floor behavior.
Figure 8 illustrates the performance curves of MBP4+ADOSD4 and AMBP4 on the (4.8.8) color codes with distances ranging from 3 to 17. The results show that MBP4+ADOSD4 using $\alpha=1.6$ improves AMBP4 starting from $d=5$.
The advantages of our BP+OSD schemes over AMBP4 are also evident in rotated toric, rotated surface, XZZX, and (6.6.6) color codes. Additionally, MBP4+OSD4-2 and MBP4+ADOSD4exhibit nearly identical performance on these codes.
Figure 8: AMBP4 vs. MBP4+OSD4-2 on the (4.8.8) color codes.
To sum up, ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes. The threshold analysis for these codes in the next subsection will reflect these results as well. We remark, however, the time consumption for ADOSD4 is significantly lower than that for OSD4-2, as will be shown in Section VI-G.
VI-F Threshold analysis for the topological codes
The decoding performance of a family of 2D topological codes is typically evaluated by the its simulated error threshold, which can be estimated using the scaling ansatz method [66, 67], which is explained as follows.
Suppose that the threshold value of a code family is $\epsilon_{\mathrm{th}}$. For a code in this family of distance $d$, if the logical error rate is $p_{L}$ at physical error rate $\epsilon$, then
$p_{L}=f(d^{\frac{1}{\nu}}(\epsilon-\epsilon_{\mathrm{th}})),$ (14)
where $f$ is a low-degree polynomial and $\nu$ is the critical exponent. The scaling ansatz method involves fitting this equation with the best values of $\epsilon_{\mathrm{th}}$, $\nu$, and $f$.
Figure 9: The threshold of BP4+OSD4-2 on the surface codes is about $17.67\%$ from the simulations. The dashed line stands for no error correction.
@@ VI-F Threshold analysis for the topological codes
Figure 10: The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach. The estimation process begins with $d=7$ and continues up to $d=17$, ensuring that no biased information influences the results.
The threshold analysis for rotated surface codes is presented in Figure 9 and Figure 10. For each data point, we collected at least 10,000 logical error events. The analysis reveals a threshold of approximately 17.67% for BP4+OSD4-2 on surface codes. Similar analyses are conducted for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes, using both BP4+OSD4-2 and BP4+ADOSD4. In the simulations, the data is fitted using a low-degree polynomial.
Note that MBP4 with $\alpha=1$ is used in the threshold analyses. Optimizing the value of $\alpha(\epsilon)$ for MBP can enhance overall performance.
The summarized results, including thresholds obtained with AMBP4, are shown in Table II. These results are consistent with the simulations in the previous subsection. Notably, BP4+OSD4-2 shows a significant enhancement over AMBP4 for surface and color codes.
@@
TABLE II: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
@@
We also compare the thresholds of our decoder with those of other known decoders in Table III, including minimum-weight perfect matching (MWPM) [68, 69]. In [43], BP2+mOSD-(2,60) achieved a threshold of $p=9.9\%$ on toric codes for bit-flip errors. We estimate that it has a threshold of $\epsilon_{\mathrm{th}}=9.9\%\times\frac{3}{2}=14.85\%$ for depolarizing errors [1].
TABLE III: Thresholds of several decoders over depolarizing errors.
[…]
: estimated from the threshold of 9.9% in the bit-flip channel [43].
@@ VI-G Time consumption
TABLE IV: Comparison of the time consumption between ADOSD4 and OSD4-2 at logical error rate about $2\times 10^{-6}$.
Code BP failure ADOSD4 OSD4-2 time LER
rate time ($\mu$s) time ($\mu$s) ratio ratio
toric 27.76% 88.63 4712.9 1.88% $\approx 0$dB
surface 67.69% 71.97 3549 2.82% $\approx 0$dB
(6.6.6) 10.32% 35.42 3812 0.93% -2.3dB
(4.8.8) 7.17% 26.50 1265 2.09% -3.9dB
XZZX 11.93% 22.29 1001.3 2.23% $\approx$ 0dB
GHP 0.86% 1188.7 45910.3 2.59% 2.49dB
: the column ‘time ratio’ represents the ratio of MBP4+ADOSD4 time to OSD4-2 time.
: the column ‘LER ratio’ represents the ratio of the logical error rate using ADOSD4 to that using OSD4-2. If the LER ratio is greater than 0dB, it means that ADOSD4 is better.
In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes.
VI-G1 Time Consumption of ADOSD4 vs. OSD4-2
First, we compare the time consumption of OSD4-2 and ADOSD4 for various topological codes of comparable distances with the [[882, 48,16]] GHP code, including the [[324,2,18]] rotated toric code, the [[289,1,17]] rotated surface code, the [[[217, 1,17]] (6.6.6) color code, the [[161, 1,17]] (4.8.8) color code, and the [[145, 1,17]] twisted XZZX code. For each code, the simulated physical error rate is chosen so that the logical error rate of MBP4+ADOSD4 is approximately $2\times 10^{-6}$. We fix $\alpha=1$ in this subsection and the decoding performance can be further improved.
In the simulations, when BP fails, the BP outputs will be handled by ADOSD4 and OSD4-2, respectively, We note that ADOSD4 is invoked only when MBP4 fails. and their time consumption statistics will be counted separately in the following tables. For example, ADOSD4 time only counts the time to run Algorithm 4 when MBP4 fails.
The simulation results are summarized in Table IV and the unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer. It can be observed that the time consumption of ADOSD4 is roughly $1\%$ to $3\%$ of the time consumption for OSD4-2 at logical error rate about $2\times 10^{-6}$ for the various quantum codes.
For the toric, surface, and twisted XZZX codes, ADOSD4 and OSD4-2 exhibit almost identical performances, while OSD4-2 performs better on the color codes and ADOSD4 performs better on the GHP code. This is reflected in the last column of Table IV.
To summarize, ADOSD4 is at least 30 times faster than OSD4-2, with a maximum performance loss of 4dB for topological codes at logical error rate $2\times 10^{-6}$. Additionally, for the [[882, 48]] GHP code, ADOSD4 also enhances performance.
VI-G2 Time Consumption of ADOSD4 vs. BP
We have demonstrated that ADOSD4 is generally faster than OSD4-2. Next, we compare the time consumption of ADOSD4 to MBP4, using surface codes.
TABLE V: Comparison of the time consumption for surface codes at logical error rate about $10^{-6}$.
code average time per MBP4 time per time per
distance iterations MBP4 failure ADOSD4 OSD4-2
iteration ($\mu$s) rate ($\mu$s) ($\mu$s)
11 22.3473 0.01464 21.46% 0.0272 0.667
13 37.1148 0.0212 35.97% 0.0469 1.342
15 52.6400 0.0287 51.47% 0.0506 2.155
: average over all samples, including MBP failures.
TABLE VI: Time Consumption of ADOSD4 vs. BP for the surface codes
code $\epsilon$ average iterations ADOSD4 time
distance when MBP4 succeeds in MBP4 iterations
11 1.7% 1.002 2
13 2.5% 1.79 2.21
15 3.3% 2.41 1.76
: average over only the samples when MBP4 succeeds.
Table V lists the time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15. To achieve a logical error rate $10^{-6}$ using ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively. At least $10^{5}$ samples are collected for each case. For example, at $d=11$, the MBP4 failure rate is $21.46\%$ at physical error rate 1.7%. Since the maximum number of iterations is set to be $T=100$, this contributes to an average of $21.46$ iterations. Since the average iterations include MBP4 failures is 22.3473, the average iterations when MBP4
Table VI summarizes the time consumption of ADOSD4 in terms of MBP4 iterations. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations. This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions. Furthermore, the time usage for ADOSD4 is significantly smaller than OSD4-2.
VI-G3 HRSR and OSD-0 in ADOSD4
@@
Table VII enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10. At $\epsilon=4.1\%$, 93.61% of ADOSD4 trials already perform only OSD4-0, increasing to 99.84% at $\epsilon=1.7\%$. Thus, at lower error rates, almost all trials execute OSD4-0.
Moreover, the HRSR algorithm effectively reduces the dimensions of the check matrix in both rows and columns. Table VII also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table). These percentages increase quickly as the error rate decreases. With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the results in Table IV.
TABLE VII: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
@@ VII Conclusion
Although the intention of [56] and [57] is quite similar to our HRSR algorithm, the underlying approaches differ: their methods focus on dividing the decoding problem into several smaller problems, whereas our approach involves simply removing the most reliable bits. Consequently, it may be possible to combine their methods with ours. For example, apply HRSR first, followed by either of the algorithms.
@@
An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Lemma 8 and Corollary 10. This information can be used to improve the algorithm by flipping effective bits in approximate degenerate decoding. Additionally, another potential improvement is extending the candidate pruning condition by Fossorier et al[46] to quantum codes.
@@
There are various reliability metrics that consider the (weighted) accumulation of the log-likelihood ratio (LLR) signs for each bit during BP iterations [70, 71, 72, 73]. Our hard-decision reliability can be viewed as an extension of this idea, focusing on the length of the last run in the hard-decision vectors. The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max. From our experiments, the integrated hard- and soft-decision reliabilities yielded the best results.
Finally, our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [74, 75], using techniques proposed in [76, 77, 78], or to the circuit-level noise model [79], as demonstrated in our recent work [78].
Currently, there is significant progress in circuit-level error decoding [80, 81, 55, 82]. Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS [43], guided decimal guessing (GDG)[54], ambiguity clustering[56], localized statistics decoding (LSD)[57], and ordered Tanner forest (OTF) [83]. These decoders are based on detector error models generated by STIM [84] and utilize binary BP combined with postprocessing.
However, there are several code capacity error decoding problems where BP-OSD is suboptimal, indicating potential for further improvement in the circuit-level noise model. Therefore, we aim to extend our BP+OSD method to address these more demanding scenarios effectively.

### v2

@@ VI-B Robustness of the maximum iterations in MBP4+ADOSD4
Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations ${\mathrm{T}}$ for the $[[175,7,7]]$ LCS code.
Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. We note that in ADOSD4, OSD-$w$ is invoked when HRSR fails. As a result, the overall algorithm maintains high accuracy. However, the failures of the HRSR algorithm can impact the simulation time, as OSD-$w$ is significantly more time-consuming.
Figure 3 illustrates the performance of MBP4+ADOSD4 decoding for the [[175,7,7]] LCS code with varying maximum numbers of iterations, ${\mathrm{T}}=1,2,3,4,5,100,200$. The curve for ${\mathrm{T}}=4$ is already good enough and smooth. While increasing the number of iterations results in slightly improved performance, the curves converge closely around an LER of $2\times 10^{-7}$ to $3\times 10^{-7}$.
$\epsilon$ 0.1 0.056 0.032 0.018 0.01 0.007 0.0056
Avg. iter. 194.46 97.67 23.51 6.31 2.44 1.70 1.23
TABLE II: The average number of iterations for MBP4 decoding on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=200$.
Table II presents the average number of iterations for ${\mathrm{T}}=200$. It can be observed that the average number of iterations rapidly decreases to one or two when $\epsilon<0.01$. Since we are targeting lower error rates, to balance accuracy and efficiency, we typically choose ${\mathrm{T}}=100$.
VI-C Selection of Reliability Threshold
We now discuss the selection of the reliability threshold. Figure 4 presents an example for the $[[175,7,7]]$ LCS code, which is representative of our simulation results.
Based on our experience, at high error rates, the reliability threshold has little impact on performance. However, at low error rates, where higher accuracy is required, the choice of $\theta$ depends on both ${\mathrm{T}}$ and the specific code. For instance, MBP4+ADOSD4 achieves excellent performance on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=2$ and $\theta=0.5$ but performs poorly on the (4.8.8) color code with $d=5$ under the same setting.
With ${\mathrm{T}}=100$ and $\theta=0.999995$, we generally observe good and stable decoding performance across all codes considered, as shown in Figure 4. Therefore, we adopt these values for our analysis in the following subsections.
Figure 4: MBP4+ADOSD4 decoding with various values of reliability threshold $\theta$ for the $[[175,7,7]]$ LCS code.
@@ VI-D Generalized hypergraph-product codes
The $[[882,48,16]]$ GHP code [43] is highly degenerate, which causes conventional BP to perform poorly. Figure 5 presents the performance curves of different MBP and OSD combinations, where $\alpha$ is either fixed at 1.6 or adaptively chosen as $\alpha(\epsilon)$.
For comparison, we also include the performance of AMBP4 [32] and BP+mOSD($w$, $\lambda$[43]. BP+mOSD($w$, $\lambda$) exhibits an early error floor around $\epsilon=0.09$. Both MBP4+ADOSD4 and MBP4+OSD4-2 with $\alpha=1.6$ outperform BP+mOSD($15$, $15$[43] by more than an order of magnitude at $\epsilon=0.05$.
[…]
Figure 5: Comparison of various decoders on the $[[882,48,16]]$ GHP code. The curve BP+mOSD(15,15) is from [43].
Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code with $\alpha=1.6$.
In ADOSD4, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4). Herein we also simulate the performances of MBP4+mOSD based on the reliability metrics of Entropy or Max as shown in Figure 6, including mOSD4  as well. Two conclusions are as follows:
@@
VI-E Bivariate bicycle codes
BB codes are a class of QLDPC codes proposed in [56], featuring a high error threshold and a significantly higher code rate than topological codes. Several attempts at decoding this code family have been presented in [54, 55] under the code capacity noise model. However, the BP+CB decoder in [54] does not perform as well as BP-OSD-0. The BP+GDG decoder in [55] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [55, Figure 4].
For reference, we include dotted $O(\epsilon^{t+1})$ curves for each code. The $[[72,12,6]]$ and $[[90,8,10]]$ codes closely follow these reference curves, while the other three codes exhibit lower error floors below LER $10^{-7}$.
VI-F 2D topological codes
Figure 8: MBP4+OSD4-2, MBP4+ADOSD4, and AMBP4 decoding performance for various $[[d^{2},1,d]]$ rotated surface codes.
@@ VI-E Bivariate bicycle codes
Figure 9: (a) The threshold of MBP4+OSD4-2 on the $[[d^{2},1,d]]$ rotated surface codes with $\alpha=1.0$ is about $17.67\%$ from the simulations. For each data point, we collected at least 10,000 logical error events. The dashed line stands for no error correction. (b) The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach. The estimation is based on data from $d=7$ to $d=17$, ensuring that the results are not influenced by biased information.
Here, we evaluate the decoding performance of our BP+OSD schemes on various 2D topological codes. Examining LER around $10^{-6}$ is important, as it helps identify whether an iterative decoder exhibits error floor behavior.
Figure 8 presents the performance curves of MBP4+ADOSD4 and MBP4+OSD4-2 on $[[d^{2},1,d]]$ rotated surface codes with distances ranging from 5 to 17. They exhibit excellent performance, closely following the dotted reference curves.
The threshold analysis for $[[d^{2},1,d]]$ rotated surface codes, conducted using the scaling ansatz method [67, 68], is shown in Figure 9. The results show a threshold of approximately 17.67% for BP4+OSD4-2, with MBP4+ADOSD4 exhibiting nearly identical performance. Notably, they outperform AMBP4, particularly for larger $d$, resulting in a higher threshold compared to the  16% threshold of AMBP4.
In summary, our decoding schemes demonstrate superior performance on surface codes compared to the MWPM decoder [69, 70], both in terms of LER performance and error threshold (see, e.g., [71]).
Similar analyses are performed for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes using both BP4+OSD4-2 and BP4+ADOSD4. The summarized threshold results are presented in Table III. ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes.
Figure 10: AMBP4 and MBP4+ADOSD4 decoding performance for various (4.8.8) color codes.
@@
TABLE III: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
@@
The thresholds for the color codes are lower than those of toric and surface codes. This is evident from the performance curves of the (4.8.8) color codes as shown in Figure 10. While the slopes of these curves increase as the error rate decreases, they do not closely match the slopes of the dotted reference curves. This suggests that some low-weight errors, even those with weight less than $t+1$, may remain uncorrected by these decoders. Consequently, additional techniques or higher-order OSD may be necessary.
TABLE IV: Comparison of the time consumption for surface codes at LER about $10^{-6}$.
code $\epsilon$ MBP4 time per MBP4 time per time per average iterations ADOSD4 time
distance failure rate iteration ($\mu$s) ADOSD4 ($\mu$s) OSD4-2 ($\mu$s) when MBP4 succeeds in MBP4 iterations
11 1.7% 21.46% 0.01464 0.0272 0.667 1.002 2
13 2.5% 35.97% 0.0212 0.0469 1.342 1.79 2.21
15 3.3% 51.47% 0.0287 0.0506 2.155 2.41 1.76
${}^{\dagger}:$ the time per ADOSD4 operation divided by the time per MBP4 iteration.
@@ VI-G Time consumption
In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes. We fix $\alpha=1$ in MBP4.
VI-G1 Time Consumption of ADOSD4 vs. OSD4-2 and BP
Table IV lists the average time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15 at LER arouond $10^{-6}$. The unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer. To achieve LER $10^{-6}$ using MBP4+ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively. At least $10^{5}$ samples are collected for each case.
It can be observed that the time consumption of ADOSD4 is approximately 2% to 4% of that of OSD4-2, with similar results observed for other quantum codes in this study. In summary, ADOSD4 is at least 25 times faster than OSD4-2 at an LER of around $10^{-6}$.
Table IV also lists the time consumption of ADOSD4 in terms of MBP4 iterations. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations. This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
VI-G2 HRSR and OSD-0 in ADOSD4
@@
Table V enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10. At $\epsilon=4.1\%$, 93.61% of ADOSD4 trials already perform only OSD4-0, increasing to 99.84% at $\epsilon=1.7\%$. Thus, at lower error rates, almost all trials execute OSD4-0.
Moreover, the HRSR algorithm reduces the dimensions of the check matrix in both rows and columns. Table V also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table). These percentages increase quickly as the error rate decreases. With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the previous results.
TABLE V: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
@@ VII Conclusion
An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Theorem 8 and Corollary 10. This information can be used to improve the algorithm by flipping effective bits in approximate degenerate decoding. Additionally, another potential improvement is extending the candidate pruning condition by Fossorier et al[47] to quantum codes.
@@
There are various reliability metrics that consider the (weighted) accumulation of the log-likelihood ratio (LLR) signs for each bit during BP iterations [72, 73, 74, 75]. Our hard-decision reliability can be viewed as an extension of this idea, focusing on the length of the last run in the hard-decision vectors. The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max [43]. From our experiments, the integrated hard- and soft-decision reliabilities yielded the best results.
Finally, there is significant progress in circuit-level error decoding [76, 77, 78, 56, 79]. Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS, GDG, ambiguity clustering[57], LSD [58], and ordered Tanner forest [80]. Our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [81, 82], using techniques proposed in [83, 84, 85], as demonstrated in our recent work [86].